OpenStudy (shallowgam3r):

Find the x-intercepts of the parabola with vertex (-3,16) and y-intercept (0,-20). Write your answer in this form: (x,y),(x,y). If necessary, round to the nearest hundredth.

OpenStudy (shallowgam3r):

someone please help

OpenStudy (shallowgam3r):

i only have about 14 minutes to answer this question..

OpenStudy (shallowgam3r):

@freckles @brill i keep seeing these two names. can one of you please help me

OpenStudy (anonymous):

I would recommend using vertex form Which is \[y=a(x-h)^2+k\]

OpenStudy (anonymous):

This is where h and k are the vertex coordinates in the form (h,k) In this case the vertex, (-3, 16), would be substituted as \[y=a(x+3)^2-16\]

OpenStudy (anonymous):

You can now substitute other point in, for example the y-intercept is (0,-20) so you can substitute that in for x and y and solve So that would be \[-20=a(0+3)^2-16\] What would a be when you solve?

OpenStudy (anonymous):

5/7=a/35